Three resistors connected in parallel have individ
Answer #1

The combined resistance is given as

R = 2 + \left ( \frac{1}{4}+\frac{1}{6} +\frac{1}{10}\right )^{-1}

R = 3.935 \Omega

Current in the circuit is

I = \frac{V}{R} = 12/3.935 = 3.05

The voltage across the parallel combinations is

V' = 12 - (2) 3.05 = 5.9 volts

Current in 10 ohms resistor is

I_{10} = \frac{5.9 volts}{10} = 0.59 A